\begin{thebibliography}{LKM16}

\bibitem[BL09]{BandiniLonghi}
Andrea Bandini and Ignazio Longhi.
\newblock Selmer groups for elliptic curves in {$\Bbb Z^d_l$}-extensions of function fields of characteristic {$p$}.
\newblock {\em Ann. Inst. Fourier (Grenoble)}, 59(6):2301--2327, 2009.

\bibitem[CS05]{coatessujatha05}
J.~Coates and R.~Sujatha.
\newblock Fine {S}elmer groups of elliptic curves over {$p$}-adic {L}ie extensions.
\newblock {\em Math. Ann.}, 331(4):809--839, 2005.

\bibitem[FW79]{ferrerowash}
Bruce Ferrero and Lawrence~C. Washington.
\newblock The {I}wasawa invariant {$\mu \sb{p}$}\ vanishes for abelian number fields.
\newblock {\em Ann. of Math. (2)}, 109(2):377--395, 1979.

\bibitem[GJS]{GhoshJhaShekhar}
Sohan Ghosh, Somnath Jha, and Sudhanshu Shekhar.
\newblock {I}wasawa theory of fine {S}elmer groups over global fields.
\newblock {\em Arxiv: 2201.01751}.

\bibitem[Gre73]{Greenberg73}
Ralph Greenberg.
\newblock The {I}wasawa invariants of {${\bf \Gamma }$}-extensions of a fixed number field.
\newblock {\em Amer. J. Math.}, 95:204--214, 1973.

\bibitem[Iwa59]{IwasawaMain}
Kenkichi Iwasawa.
\newblock On {$\Gamma $}-extensions of algebraic number fields.
\newblock {\em Bull. Amer. Math. Soc.}, 65:183--226, 1959.

\bibitem[JS11]{jhasujatha}
Somnath Jha and R.~Sujatha.
\newblock On the {H}ida deformations of fine {S}elmer groups.
\newblock {\em J. Algebra}, 338:180--196, 2011.

\bibitem[Kle17]{Kleine}
S\"{o}ren Kleine.
\newblock Local behavior of {I}wasawa's invariants.
\newblock {\em Int. J. Number Theory}, 13(4):1013--1036, 2017.

\bibitem[Kle21]{KleineCanad}
S\"{o}ren Kleine.
\newblock Bounding the {I}wasawa invariants of {S}elmer groups.
\newblock {\em Canad. J. Math.}, 73(5):1390--1422, 2021.

\bibitem[KT03]{KatoTrihan}
Kazuya Kato and Fabien Trihan.
\newblock On the conjectures of {B}irch and {S}winnerton-{D}yer in characteristic {$p>0$}.
\newblock {\em Invent. Math.}, 153(3):537--592, 2003.

\bibitem[LKM16]{LimMurty}
Meng~Fai Lim and V.~Kumar~Murty.
\newblock The growth of fine {S}elmer groups.
\newblock {\em J. Ramanujan Math. Soc.}, 31(1):79--94, 2016.

\bibitem[Mil86]{Milne1986}
James~S. Milne.
\newblock {\em Arithmetic duality theorems}, volume~1 of {\em Perspectives in Mathematics}.
\newblock Academic Press, Inc., Boston, MA, 1986.

\bibitem[Ros02]{Rosen}
Michael Rosen.
\newblock {\em Number theory in function fields}, volume 210 of {\em Graduate Texts in Mathematics}.
\newblock Springer-Verlag, New York, 2002.

\bibitem[She18]{shekhar}
Sudhanshu Shekhar.
\newblock Comparing the corank of fine {S}elmer group and {S}elmer group of elliptic curves.
\newblock {\em J. Ramanujan Math. Soc.}, 33(2):205--217, 2018.

\bibitem[Tan10]{Tan}
Ki-Seng Tan.
\newblock A generalized {M}azur's theorem and its applications.
\newblock {\em Trans. Amer. Math. Soc.}, 362(8):4433--4450, 2010.

\bibitem[Was97]{washingtonbook}
Lawrence~C. Washington.
\newblock {\em Introduction to cyclotomic fields}, volume~83 of {\em Graduate Texts in Mathematics}.
\newblock Springer-Verlag, New York, second edition, 1997.

\bibitem[Wut07]{wuthrichfine}
Christian Wuthrich.
\newblock Iwasawa theory of the fine {S}elmer group.
\newblock {\em J. Algebraic Geom.}, 16(1):83--108, 2007.

\end{thebibliography}

\documentclass[12pt, reqno]{amsart}
\usepackage[left=1in, right=1in]{geometry} 
\usepackage{hyperref}
\usepackage[utf8]{inputenc}
\usepackage{amssymb}
\usepackage{amscd,amsfonts,amsmath,amssymb, bbm,dsfont, comment}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}

\usepackage{enumerate,epsf,fancyhdr,float,graphicx,tabularx}
\usepackage{latexsym,mathrsfs,multirow}
\usepackage{wasysym}
\usepackage{xypic}
\usepackage[all]{xy}\usepackage[OT2,T1]{fontenc}
\usepackage[modulo,mathlines,displaymath, running]{lineno}
\usepackage{amsmath,mathrsfs,enumerate,wasysym}
\usepackage{xypic,hhline}
\usepackage[all]{xy}
\usepackage[OT2,T1]{fontenc}
\usepackage[modulo,mathlines,displaymath]{lineno}
\usepackage{tikz,tikz-cd}
\usepackage{dashrule}
\usetikzlibrary{matrix}
\usepackage{ tipa }
\usepackage[modulo,mathlines,displaymath]{lineno}

\newtheorem*{conjecture*}{Conjecture}

\newtheorem{theorem}{Theorem}[section]

\DeclareSymbolFont{cyrletters}{OT2}{wncyr}{m}{n}\DeclareMathSymbol{\Sha}{\mathalpha}{cyrletters}{"58}
\DeclareMathSymbol{\FSha}{\mathalpha}{cyrletters}{"11}
\renewcommand{\phi}{{\varphi}}

\renewcommand{\geq}{\geqslant}
\renewcommand{\leq}{\leqslant}

\newcommand{\cyc}{\mathrm{cyc}}

\newcommand{\links}{\left(\begin{array}{cc}}
\newcommand{\rechts}{\end{array}\right)}
\newcommand{\bai}{\left[\begin{array}{cc}}
\newcommand{\dai}{\end{array}\right]}
\newcommand{\hidari}{\left(\begin{array}{c}}
\newcommand{\migi}{\end{array}\right)}

\newcommand{\Ff}{\mathbb{F}}

\newcommand{\N}{\mathbb{N}}

\newcommand{\Q}{\mathbb{Q}}

\newcommand{\Z}{\mathbb{Z}}

\newcommand{\calE}{\mathcal{E}}

\newcommand{\calO}{\mathcal{O}}

\newcommand{\calU}{\mathcal{U}}
\newcommand{\calV}{\mathcal{V}}

\newcommand{\gal}{\mathrm{Gal}}

\newcommand{\Gal}{\operatorname{Gal}}

\newcommand{\fla}{\operatorname{fl}}

\newcommand{\pr}{\operatorname{pr}}

\newcommand{\coker}{\operatorname{coker}}

\newcommand{\injects}{\hookrightarrow}
\newcommand{\lra}{\longrightarrow}

\renewcommand{\contentsname}{Contents\\{\footnotesize\normalfont(A table
of contents should normally not be included)}}

\newtheorem{auxiliary proposition}[theorem]{Auxiliary Proposition}

\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{definition}[theorem]{Definition}

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{main conjecture}[theorem]{Main Conjecture}
\newtheorem{main theorem}[theorem]{Main Theorem}
\newtheorem{modesty proposition}[theorem]{Modesty Proposition}

\newtheorem{open problem}[theorem]{Open Problem}

\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{remark}[theorem]{Remark}

\newtheorem{convergence lemma}[theorem]{Convergence Lemma}
\newtheorem{corrected lemma}[theorem]{Corrected Lemma}
\newtheorem{growth lemma}[theorem]{Growth Lemma}
\newtheorem{coefficient lemma}[theorem]{Integrality Lemma}
\newtheorem{interpolation lemma}[theorem]{Interpolation Lemma}
\newtheorem{kernel lemma}[theorem]{Kernel Lemma}
\newtheorem{limit lemma}[theorem]{Limit Lemma}
\newtheorem{tandem lemma}[theorem]{Modesty Lemma}
\newtheorem{zero-finding lemma}[theorem]{Zero-Finding Lemma}

\usepackage{hyperref}

\hypersetup{
	colorlinks=true,
	linkcolor=black,
	filecolor=magenta,      
	urlcolor=black,
	pdftitle={Overleaf Example},
	pdfpagemode=FullScreen,
}

\title[~]{ A Study of Fine Selmer Groups Over Function Fields via Greenberg Neighbourhoods}
\author[S. Ghosh]{Sohan Ghosh}
\address[Ghosh]{Indian Institute of Science Education and Research, Mohali,Punjab (140306) INDIA}
\email{ghoshsohan4@gmail.com}

\keywords{Iwasawa theory; fine Selmer groups; Elliptic Curves; function field}
\subjclass[2020]{Primary: 11R23,  Secondary: 20C07, 11G05, 11R58  }
\begin{document}

\maketitle
\begin{abstract}
In \cite{Greenberg73}, Greenberg examined the local behavior of Iwasawa invariants as functions on the set $\mathcal{E}(F)$, the set of all $\mathbb{Z}_p$-extensions of a number field $F$. Kleine \cite{KleineCanad} later extended these ideas to explore the variation of Iwasawa invariants in the context of Selmer groups of elliptic curves across different $\mathbb{Z}_p$-extensions of $F$.

Let $K$ be a global function field of characteristic $p$.
In this article, we investigate the relation between Iwasawa invariants of fine Selmer groups of an elliptic curve over $K$ across various $\mathbb{Z}_p$-extensions of $K$, utilizing Kleine's techniques. Furthermore, we connect this analysis to an analogue of  \nameref{conja} by Coates and Sujatha \cite{coatessujatha05} for different $\mathbb{Z}_p$-extensions of the function field $K$.
\end{abstract}
\section{Introduction}
\noindent Let us fix a prime $p$ throughout. Iwasawa theory originated from the work of K. Iwasawa, who studied the growth of the $p$-primary parts of class groups in towers of $\mathbb{Z}_p$-extensions. He established the following celebrated theorem:

\begin{theorem}[Iwasawa]\cite{IwasawaMain} 
Let $F$ be a number field and $F_\infty$ be a $\mathbb{Z}_p$-extension of $F$. For each $n\geq 0$, let $p^{e_n}$ denote the exact power of $p$ that divides $\#Cl(F_n)$. Then there exist invariants $\lambda, \mu \in \mathbb{Z}_{\geq 0}$ and $\nu \in \mathbb{Z}$, depending on $p$ but independent of $n$, such that $e_n = \lambda n + \mu p^n + \nu$ for $n \gg 0$. \end{theorem}

\noindent For the cyclotomic $\mathbb{Z}_p$-extension $F_\cyc$ of $F$, Iwasawa further conjectured that $\mu = 0$. Ferrero and Washington \cite{ferrerowash} later proved this conjecture for abelian number fields, but it remains open for general number fields.

In 2005, Coates and Sujatha \cite{coatessujatha05} found a surprising connection between Iwasawa's $\mu=0$ conjecture and the structure of fine Selmer groups of elliptic curves. More precisely, they proved the following important theorem:

\begin{theorem}\cite[Theorem 3.4]{coatessujatha05}\label{thm
} Let $E/F$ be an elliptic curve, and let $p$ be an odd prime such that $F(E_{p^\infty})/F$ is a pro-$p$ extension. Then $R(E/F_\text{cyc})^\vee$ is a finitely generated $\mathbb{Z}_p$-module if and only if Iwasawa’s   $\mu = 0$ conjecture holds for $F_\cyc$. \end{theorem}

\par Motivated by this result, Coates and Sujatha formulated the following conjecture:

\begin{conjecture*}[Conjecture A]\label{conja} 

For all elliptic curves $E$ over $F$, $R(E/F_\text{cyc})^\vee$ is a finitely generated $\mathbb{Z}_p$-module. 

\end{conjecture*}

\par Following   \nameref{conja},  the fine Selmer groups over  cyclotomic $\Z_p$-extensions of a number field have been studied  by various authors including \cite{LimMurty,jhasujatha, shekhar, wuthrichfine} etc.

A natural question  is how the Iwasawa invariants vary across different $\mathbb{Z}_p$-extensions of $F$. Greenberg addressed this by introducing a topology on $\mathcal{E}(F)$, the set of all $\mathbb{Z}_p$-extensions of $F$, allowing for a comparative analysis of Iwasawa invariants across distinct $\Z_p$-extensions. Using Fukuda modules, Kleine \cite{Kleine} examined the local behavior of Iwasawa invariants across $\mathcal{E}(F)$. He later generalized these ideas to Selmer groups associated with elliptic curves having good-ordinary reduction at primes above $p$ \cite{KleineCanad}.

In \cite{GhoshJhaShekhar}, the authors initiated the study of fine Selmer groups over function fields of characteristic $p$. These fine Selmer groups are subgroup of the classical Selmer groups. In this article, we investigate the variation of Iwasawa invariants of fine Selmer groups over certain neighborhoods of $\mathbb{Z}_p$-extensions of a function field $K$, which is defined over a finite field of characteristic $p$. Our main results are Theorem \ref{mainthm:1} and Theorem \ref{mainthm:2}. We are also working on generalizing these ideas to Selmer groups in both equal and unequal characteristic case over function fields as a part of our future project.

The structure of the article is as follows: In \S \ref{Preliminaries}, we review key definitions and preliminaries necessary for the remainder of the article. In \S \ref{mainresults}, we present the proofs of our main results. Finally, in \S \ref{example}, we conclude with a discussion of an example.

%Understanding these variations is crucial not only for the study of Iwasawa theory in number fields but also for its analogs in function fields, where many phenomena from classical number theory find new and intriguing parallels. 

\section{Preliminaries}\label{Preliminaries}
\par Let $C_K$ be a curve over the finite field $\mathbb{F}$ of characteristic $p$, with $K$ as its function field. Let $E/K$ be an elliptic curve. Consider $U$ to be a dense open subset of $C_K$ such that $E/K$ has good reduction at every place in $U$.  Denote by $\Sigma_K$ the set of all places of $K$, and let $S$ represent the set of places of $K$ outside $U$. 

Let $L$ be a finite extension   of $K$ inside $K_S$, the maximal algebraic extension of $K$ unramified outside $S$. Let $v$ be any  prime of $K$ and $w  $ denote  a prime of $L$. Define

\begin{equation}\label{equationdefine}
	\begin{aligned}
		J_v^1(E/L):=\underset{w\mid v}\bigoplus	\frac{H^1_{\fla}(L_w, E_{p^\infty})}{im(\kappa_w)} \ \text{and }
		K_v^1(E/L):= \underset{w\mid v}\bigoplus	H^1_{\fla}(L_w, E_{p^\infty}).	
	\end{aligned}
\end{equation}

 Here $H^i_{\fla}(-,-) $ denote the flat cohomology \cite[Chapters II, III]{Milne1986} and $\kappa_w: E(L_w)\otimes {\Q_p}/{\Z_p}\injects H^1_{\fla}(L_w,E_{p^\infty})$ is induced by the Kummer map. The following definition of the Selmer group is recalled from \cite{KatoTrihan}:
\begin{definition}\cite[Prop. 2.4]{KatoTrihan}
	With $\Sigma_K, S \text{ and } K\subset L \subset K_S$ as above, define 
	\begin{small}
	\begin{equation}\label{definition24}
		S(E/L):= \ker\big(H^1_{\fla}(L, {E}_{p^\infty})\lra \underset{v\in \Sigma_K}\bigoplus J_v^1(E/L)\big).
	\end{equation}
\end{small}
	We define the $S$-fine Selmer group as:
	\begin{small}
	\begin{equation}\label{definition4.1}
		\begin{aligned}
			R^S(E/L):={} \ker\left(H^1_{\fla}(L, E_{p^\infty})\lra \underset{v\in S}\bigoplus K_v^1(E/L) \underset{v\in \Sigma_K\setminus S}\bigoplus J_v^1(E/L)\right)\\
			\cong \ker\left (S(E/L)\lra \underset{w\mid v, v\in S}\bigoplus E(L_w)\otimes \Q_p/\Z_p\right).
		\end{aligned}
	\end{equation}
\end{small}
\end{definition}
For an infinite algebraic extension $\mathcal{L}$ of $K$, these definitions extend as usual by taking the inductive limit over finite subextensions of $\mathcal{L}$ over $K$.

\subsection{ Examples of $\Z_p$-extensions of function field}
 Let $\Ff^{(p)}$ be the unique subfield of $\overline{\Ff}$, the algebraic closure of $\Ff$, such that $\gal(\Ff^{(p)}/\Ff)\cong \Z_p$. Set $K_\infty:=K\Ff^{(p)}$. Note that $K_\infty/K$ is unramified everywhere.  This $\Z_p$-extension $K_\infty$ is referred to as in the literature as the "arithmetic" $\Z_p$-extension. 
 
 Unlike the number fields, in case of function fields, we can construct infinitely many $\Z_p^d$-extensions of $K$ for any $d\geq 1$. 
 
 { We will now explore a few examples of $\Z_p^d$-extensions constructed from Carlitz modules. These $\Z_p^d$-extensions are constructed in the following manner:}

 	Let $K=\Ff(t)$ and $P(t)=a_nt^n+\cdots a_0 \in \Ff [t]$. We define the Carlitz polynomial $[P(t)](X)$ with coefficients in $\Ff[t]$ recursively as follows:

		$[1] (X)=X$, 
		
		$[t] (X)=X^p +tX$,
		
		$[t^n] (X)=[t]([t^{n-1}](X))$ and
		
		$[a_n t^n+\cdots+a_1 t+a_0](X)= a_n[t^n](X)+\cdots +a_1[t](X)+a_0 (X).$
		
		{ Consider a field extension $F$ of $K$}. Then $F$ can be thought of as a $\Ff[t]$-module, where the action of $\Ff[t]$  is given by the Carlitz polynomials. 
		Choose  a prime $\mathfrak P$ of $\Ff[t]$. For $n>0$, let
		$$\Lambda_{\mathfrak P^n}:=\{\lambda\in \overline{\Ff(t)}| [\mathfrak{P}^n](\lambda)=0\}.$$
		
		{Here} $K(\Lambda_{\mathfrak P ^n})/K$ is  Galois  with $\gal(K(\Lambda_{\mathfrak P ^n})/K)\cong (\Ff[t]/\mathfrak{P}^n)^\times$. Put $\widetilde{K}:=\underset{n\geq 1}\bigcup K(\Lambda_{\mathfrak{P}^n})$, then $\gal(\widetilde K /K)\cong \Z_p^\N \times (\Ff[t]/\mathfrak{P})^\times$. 

  The $\Z_p^d$-extensions obtained from $\widetilde {K}$, for $d\geq 1$ are ramified only at the prime $\mathfrak P$ and they are totally ramified at that prime \cite[Proposition 12.7]{Rosen}. The $\Z_p$-extensions, thus obtained, are examples of  ``geometric" $\Z_p$-extensions.
\subsection{Greenberg topology}
For a $\Z_p$-extension $K_\infty/K$, we define $\mathcal{P}(K_\infty)$ as the set of primes in $K$, ramified in $K_\infty$.
Let $\mathcal{E} (K)$ be the set of all  $\Z_p$-extensions of $K$, which are finitely ramified. Define
\[
\mathcal{U}(K_\infty,n):=\{K_\infty'\in \mathcal{E} (K)| [K_\infty\cap K_\infty': K]\geq p^n, \mathcal{P}(K_\infty')\subset\mathcal{P}(K_\infty) \},
\]
for some $n\geq 0$ and $K_\infty'\in \mathcal{E}(K)$. We define a topology on $\mathcal{E} (K) $ generated by these sets.

Let $E$ be an elliptic curve defined over $K$.  Define the set
\[
S_1:=\{v\in\Sigma_K\ |\ v \textrm{ ramifies in } K_\infty\}\bigcup \{v\in \Sigma_K\ |\ E \textrm{ has bad reduction at } v\}.
\]
 Note that $S_1$ is a finite set. Let $S_2$ be another finite set of primes of $K$, where $S_1 \subseteq S_2$, and for any prime $v \in S_2 \setminus S_1$, the prime $v$ splits completely in $K_\infty$.

 We consider a refinement of the above topology: for ${K}_\infty\in \mathcal{E}(K)$ and $m\in \N$, let $\calV(E, K_\infty,m)\subset \mathcal{U}(K_\infty,m)$ denote the subset of $\Z_p$-extensions $\widetilde{K}_\infty$ of $K$ in $\mathcal{U}(K_\infty,m)$ such that each prime of $S_2\setminus S_1$ (that is totally
split in $K_\infty$)  splits completely  in $\widetilde{K}_\infty$. We consider the topology on $\mathcal{E}(K)$, generated by the sets  $\calV(E, K_\infty,m)$, {i.e, the smallest topology  on $\calE(K)$ generated by these sets. This is a finer topology on $\mathcal{E}(K)$ than the topology generated by the sets 
 $\{\mathcal{U}(K_\infty,m), m \geq 0\}$. Indeed it is easy to see that if $K_\infty^\prime \in \mathcal{U}(K_\infty,m)$, then $L_\infty \in \calV(E, K_\infty',m+1) \subset \mathcal{U}(K_\infty,m)$. Recall, that a similar type of refined neighbourhoods were also considered in \cite{KleineCanad}.
}
\subsection{Fukuda Modules}
 In \cite{Kleine}, Kleine introduced a special family of Iwasawa modules $X = \varprojlim X_n$. These modules were designed to derive information about the projective limit by utilizing data from a sufficiently large number of layers $X_n$.
\begin{definition}\label{defn:Fukudamodules}
Let $1 \le d \in \mathbb{N}$. Let $R$ be the ring of formal power series over $\mathbb{Z}_p$ in $d$ variables with maximal ideal $\mathfrak{m} = (p, T_1, \ldots, T_d)$. Let $X = \varprojlim X_n$ denote the projective limit of $R$-modules $X_n$, $n \in \mathbb{N}$, each assumed to be an abelian pro-$p$ group. Further assume that $X$ is compact as an $R$-module with respect to the $\mathfrak{m}$-adic topology. Let $Y_n \subset X$ denote the kernel of the projection map $X \to X_n$ for each $n$. Let $C_1, C_2, C_3 \in \mathbb{N}$ be $p$-powers. Then $X$ is called a Fukuda $R$-module with parameters $(C_1, C_2, C_3)$ if there exists a family of compact $R$-submodules $(Z_n)_n$ of $X$ such that 
$$|\coker(\pr_n)| \le C_1, Z_{n+1} \subset \mathfrak{m} \cdot Z_n, [Y_n : (Y_n \cap Z_n)] \le C_2,$$
and
$$[Z_n : (Y_n \cap Z_n)] \le C_3$$ for each $n \in \mathbb{N}$.
\end{definition}

Consider a $\mathbb{Z}_p^d$-extension $K_\infty/K$, where $d \geq 1$. Write $K_\infty = \bigcup_{n} K_n$ with $\Gal(K_n/K) \cong (\Z/p^n\Z)^d$. Let $X = \varprojlim X_n$, where $X_n$ is the $p$-primary subgroup of the ideal class group of $K_n$. Then, under appropriate assumptions on the ramification of primes in $K_\infty/K$, one can show that $X$ is a Fukuda-$R$-module with parameters $(1,1,1)$ (see \cite[\S 3]{KleineCanad}), where $R \cong \mathbb{Z}_p[[\Gal(K_\infty/K)]]$.

Before moving further, we recall the following result from \cite{BandiniLonghi}:
\noindent \begin{lemma}\cite[Lemma 4.1]{BandiniLonghi}\label{lemma:BandiniLonghi}
    Let $\Gamma \cong \mathbb{Z}_p^d$ and $B$ a finite $p$-primary $\Gamma$-module. Then 
\[
|H^1(\Gamma, B)| \leq |B|^d \quad \text{and} \quad |H^2(\Gamma, B)| \leq |B|^{\frac{d(d-1)}{2}}.
\]
\end{lemma}
\section{Main Results}\label{mainresults}
\par Let $K$ be a function field over the finite field of characteristic $p$. Let $K_\infty$ be a $\Z_p$-extension of $K$ that is finitely ramified and $\Gamma:=\Gal(K_\infty/K)$. Let $E$ be a non-isotrivial elliptic curve (i.e; $j(E)\notin \Ff$), with either good ordinary or split multiplicative reduction at all the primes of $K$. 

Recall the definition of $S_1$ as a finite set of primes of $K$, consisting only of those primes that are ramified in $K_\infty$ and those where $E$ has bad reduction. Similarly, recall that $S_2$ is another finite set of primes of $K$, with $S_1 \subseteq S_2$, where any prime $v \in S_2 \setminus S_1$ splits completely in $K_\infty$.

We first prove a control theorem from $R^S(E/K_n)\lra R^S(E/K_\infty)^{\Gamma_n}$, where the set  $S$ be either $S_1$ or $S_2$ .
\begin{theorem}\label{thm:control}
     Let $E$ be a non-isotrivial elliptic curve (i.e; $j(E)\notin \Ff$) defined over $K$. Let $S$ be either $S_1$ or $S_2$. Assume that all places in $S$, $E$ has 
     either good ordinary or split multiplicative reduction. 
    Then the kernel and cokernel of the natural map
    \[
    R^S(E/K_n)\lra R^S(E/K_\infty)^{\Gamma_n}
    \]
    are finite and bounded independently of $n$.
\end{theorem}
\begin{proof}
Consider the following commutative diagram:
% https://q.uiver.app/#q=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
\[\begin{tikzcd}
	0 & {R^S(E/K_\infty)^{\Gamma_n}} & {H^1_{\fla}(K_\infty, E[p^\infty])^{\Gamma_n}} & {\left(\underset{v\mid S}\bigoplus\ K^1_v(E/K_\infty)\ \underset{v\nmid S}\bigoplus\ J^1_v(E/K_\infty)\right)^{\Gamma_n}} \\
	0 & {R^S(E/K_n)} & {H^1_{\fla}(K_n, E[p^\infty])} & {\underset{v\mid S}\bigoplus\ K^1_v(E/K_n)\ \underset{v\nmid S}\bigoplus\ J^1_v(E/K_n)}
	\arrow[from=1-1, to=1-2]
	\arrow[from=1-2, to=1-3]
	\arrow[from=1-3, to=1-4]
	\arrow[from=2-1, to=2-2]
	\arrow["{\alpha_n}", from=2-2, to=1-2]
	\arrow[from=2-2, to=2-3]
	\arrow["{\beta_n}", from=2-3, to=1-3]
	\arrow[from=2-3, to=2-4]
	\arrow["{\gamma_n=\underset{v}\oplus \gamma_{n,v}}", from=2-4, to=1-4]
\end{tikzcd}\]
First, observe that $\ker(\beta_n) = H^1(\Gamma_n, E(K_\infty)[p^\infty])$. Since $E$ is a non-isotrivial elliptic curve, it is known that $E(K_\infty)[p^\infty]$ is finite \cite[Lemma 3.1]{BandiniLonghi}. Therefore, $\ker(\beta_n)$ is finite. In fact, by Lemma \ref{lemma:BandiniLonghi}, its size is bounded by $|E(K_\infty)[p^\infty]|$.

  %Note that if $v$ is ramified, it remains finitely split in every subextension $K_n$ of $K_\infty$.

Next, for any prime $v \mid S$, we consider three cases. The first case arises when $v$ is a bad prime. In this situation, for any prime $w \mid v$, $E(K_{\infty, w})[p^\infty]$ is finite (see proof of \cite[Theorem 4.12]{BandiniLonghi}). 

The second case is when $v$ is a good ramified prime. Let $w\mid v$ be a prime of $K_\infty$. Recall the short exact sequence, 
\[
    0\lra\widehat{E}(\mathcal{O}_{K_{\infty,w}})\lra E(K_{\infty,w})\lra \overline{E}(\Ff_{K_{\infty,w}}) \lra 0,
\]
where  $\widehat{E}$ and $\overline{E}$ be the formal group associated to $E$ and the reduction of $E$ at the prime of $K_{\infty,w}$ respectively.  Here, $\calO_{K_{\infty,w}}$ denote the ring of integers of $K_{\infty,w}$ and $\Ff_{K_{\infty,w}}$ denote the residue field at that prime.  By \cite[Lemma 2.5.1]{Tan}, $\widehat{E}(\mathcal{O}_{K_{\infty,w}})$ is a torsion free $\Z_p$-module.
 Therefore, $E(K_{\infty, w})[p^\infty] \cong E(\Ff_{K_{\infty,w}})[p^\infty]$. Since $K_\infty$ is a $\mathbb{Z}_p$-extension of $K$ which is ramified at finitely many places, there exists a subextension $K_n$ of $K_\infty$ such that the primes of $K_n$ that are ramified are totally ramified in $K_\infty$ \cite[Lemma 13.3]{washingtonbook}.  Consequently, $E(\Ff_{K_{\infty,w}})[p^\infty]$ is also finite. The third case occurs when $S=S_2$. Then there exists $v\in S$, which splits completely in $K_\infty$, then $\ker(\gamma_{n,v})=0$.
 Therefore, we conclude that $\ker(\gamma_{n,v})$ is finite and bounded for every $v \in S$ (by Lemma \ref{lemma:BandiniLonghi}).

For primes $v \notin S$, we have $\ker(\gamma_{n,v}) = 0$ \cite{Tan}. The result then follows by applying the snake lemma.
\end{proof}
\par As a corollary, we obtain the following result:
 \begin{corollary}
     Under the assumptions of Theorem \ref{thm:control}, $R^S(E/K_\infty)^\vee = \varprojlim R^S(E/K_n)^\vee$ is a Fukuda-$\Lambda$-module with parameters $(C_1, C_2, 1)$ for suitable $p$-powers $C_1$ and $C_2$.
 \end{corollary}
\begin{proof}
    The proof of this statement is similar to \cite[Corollary 4.2]{KleineCanad}.
\end{proof}
For the proof of the next proposition, we closely follow the proof of \cite[Theorem 4.5]{KleineCanad}
\begin{proposition}\label{prop:FUkudaS1}
Let us keep the hypothesis and assumption of Theorem \ref{thm:control}. Then there exist integers $m, C_1, C_2 \in \mathbb{N}$ such that $R^{S_1}(E/\widetilde{K}_\infty)^\vee$ is a Fukuda-$\Lambda$-module with bounded parameters $(C_1, C_2, 1)$ for each $\widetilde{K}_\infty \in \calU( K_\infty, m)$.
\end{proposition}
\begin{proof}
    For each $n>0$, we have the following commutative diagram:
   % https://q.uiver.app/#q=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
\[\begin{tikzcd}
	0 & {R^{S_1}(E/\widetilde{K}_\infty)^{\Gamma_n}} & {H^1_{\fla}(\widetilde{K}_\infty, E[p^\infty])^{\Gamma_n}} & {\left(\underset{v\mid {S_1}}\bigoplus\ K^1_v(E/\widetilde{K}_\infty)\ \underset{v\nmid {S_1}}\bigoplus J^1_v(E/\widetilde{K}_\infty)\right)^{\Gamma_n}} \\
	0 & {R^{S_1}(E/\widetilde{K}_n)} & {H^1_{\fla}(\widetilde{K}_n, E[p^\infty])} & {\underset{v\mid {S_1}}\bigoplus\ K^1_v(E/\widetilde{K}_n)\ \underset{v\nmid {S_1}}\bigoplus\ J^1_v(E/\widetilde{K}_n)}
	\arrow[from=1-1, to=1-2]
	\arrow[from=1-2, to=1-3]
	\arrow[from=1-3, to=1-4]
	\arrow[from=2-1, to=2-2]
	\arrow["{\alpha_n^{\widetilde{K}_\infty}}", from=2-2, to=1-2]
	\arrow[from=2-2, to=2-3]
	\arrow["{{\beta}_n^{\widetilde{K}_\infty}}", from=2-3, to=1-3]
	\arrow[from=2-3, to=2-4]
	\arrow["{{\gamma}_n^{\widetilde{K}_\infty}=\underset{v}\oplus \ {\gamma}_{n,v}^{\widetilde{K}_\infty}}", from=2-4, to=1-4]
\end{tikzcd}\]
Here $\Gamma_n=\gal(\widetilde{K}_\infty/\widetilde{K}_n)$. By arguments similar to \cite[Theorem 4.5]{KleineCanad}, it suffices to find an uniform bound for the kernel and cokernel of  the restriction map $\alpha_n^{\widetilde{K}_\infty}$ for every $\widetilde{K}_\infty$ in some open neighbourhood $\mathcal{U}(K_\infty,m)$ of $K_\infty$. This is equivalent to showing that the kernel and cokernel of the map $\beta_n^{\widetilde{K}_\infty}$ and the kernel of the map $\gamma_n^{\widetilde{K}_\infty}$ are bounded  for every $\widetilde{K}_\infty\in \mathcal{U}(K_\infty,m)$ of $K_\infty$, by snake lemma.

Let $v\in S_1$, such that $E$ has good ordinary reduction at $v$. Since, $K_\infty$ is ramified at $v$, we can choose  $m_0>0$  large enough such that all primes of $K$ ramifying in $K_\infty$  are totally ramified in $K_\infty/K_{m_0}$, then the sets of primes of
$K$ that ramify in $K_\infty$ and  in $\widetilde{K}_\infty\in \calU( K_\infty,m_0 + 1)$, coincide. Hence, the residue 
fields coincide as well. Now,  choose $w\mid v$ to be a prime of $\widetilde{K}_\infty$. As explained in the proof of Theorem \ref{thm:control}, we know that $\ker{\gamma}_{n,v}^{\widetilde{K}_\infty}$ is bounded by $|E(\Ff_{\widetilde{K}_{\infty,w}})[p^\infty]|$, where $\Ff_{\widetilde{K}_{\infty,w}}$ is the residue field of $\widetilde{K}_{\infty,w}$.
Therefore, for $v\in S$, where $E$ has good reduction at $v$, $\ker ({\gamma}_{n}^{\widetilde{K}_\infty})$ are finite and bounded, for every $\widetilde{K}_\infty\in \calU(K_\infty,m_0+1)$ by $|E(\Ff_{K_{\infty,w}})[p^\infty]|$.

Now, we make the following observations:
\begin{enumerate}
    \item  $E(K_\infty)[p^\infty]$ is finite as $E$ is non-isotrivial \cite[Lemma 4.3]{BandiniLonghi}.
    \item for $v\in S$, such that $E$ has split multiplicative reduction at $v$, $w \mid v$, $E(K_{\infty, w})[p^\infty]$ is finite (see proof of \cite[Theorem 4.12]{BandiniLonghi}).
    \item for $v\nmid S$, $\ker({\gamma}_{n,v}^{\widetilde{K}_\infty})=0$.
\end{enumerate}
Now, by arguments similar to \cite[Theorem 4.5]{KleineCanad} (also see \cite[Theorem 5.13]{KleineCanad}), we obtain a neighbourbood $\calU(K_\infty,r)$, such that $r> m_0$ where $\ker(\beta_n^{\widetilde{K}_\infty})$ and  $\ker({\gamma}_{n}^{\widetilde{K}_\infty})$ are finite and bounded by common constants or bounds, independent of $\widetilde{K}_\infty$, for every $\widetilde{K}_\infty\in \calU(K_\infty,r)$. This completes our proof.
\end{proof}

By arguments similar to Proposition \ref{prop:FUkudaS1}, we prove the following proposition:

\begin{proposition}\label{prop:FUkudaS1}
Let us keep the hypothesis and assumption of Theorem \ref{thm:control}. Then there exist integers $m, D_1, D_2 \in \mathbb{N}$ such that $R^{S_2}(E/\widetilde{K}_\infty)^\vee$ is a Fukuda-$\Lambda$-module with bounded parameters $(D_1, D_2, 1)$ for each $\widetilde{K}_\infty \in \calV(E, K_\infty,m)$.
\end{proposition}

Now,  mimicking  the ideas of \cite[Theorem 4.11]{KleineCanad} we obtain the following results, using Proposition \ref{prop:FUkudaS1}:
\begin{theorem}\label{mainthm:1}
   Let $K_\infty$ be a $\Z_p$-extension of $K$ that is finitely ramified. Let ${\Sigma_1}$ be a finite set of primes containing $S_1$.  Let $E$ be a non-isotrivial elliptic curve defined over $K$, with good ordinary and split multiplicative reduction at all primes of $S_1$. 

     Suppose, that $R^{S_1}(E/K_\infty)^\vee$ is a finitely generated torsion $\Lambda$-module.
    Then there exists a neighbourhood $U = U(K_{\infty}, m)$ of $K_{\infty}$ such that
\begin{itemize}
    \item $R^{\Sigma_1}(E/\widetilde{K}_{\infty})$ is a torsion $\Lambda$-module for each $\widetilde{K}_{\infty} \in U$;
    \item $\mu(R^{\Sigma_1}(E/\widetilde{K}_{\infty})) \leq \mu(R^{S_1}(E/K_{\infty}))$ for each $\widetilde{K}_{\infty} \in U$;
    \item $\lambda(R^{\Sigma_1}(E/\widetilde{K}_{\infty})) \leq \lambda(R^{S_1}(E/K_{\infty}))$ for each $\widetilde{K}_{\infty} \in U$ such that $\mu(R^{S_1}(E/\widetilde{K}_{\infty})) = \mu(R^{S_1}(E/K_{\infty}))$.
\end{itemize}
\end{theorem}
\begin{proof}
    The proof follows for $R^{S_1}(E/K_\infty)$  by argument similar to \cite[Theorem 4.11]{KleineCanad}, by application of Proposition \ref{prop:FUkudaS1}. Then we use the fact that  $R^{\Sigma_1}(E/K_\infty)\subset R^{S_1}(E/K_\infty)$.
\end{proof}
By similar arguments as in Theorem \ref{mainthm:1}, we get
\begin{theorem}\label{mainthm:2}
   Let $K_\infty$ be a $\Z_p$-extension of $K$ that is finitely ramified. Let $E$ be a non-isotrivial elliptic curve defined over $K$, with good ordinary and split multiplicative reduction at all primes of $S_2$.
   Let ${\Sigma_2}$ be a finite set of primes containing $S_2$.

     Suppose, that $R^{S_2}(E/K_\infty)^\vee$ is a finitely generated torsion $\Lambda$-module.
    Then there exists a neighbourhood $U = \calV(E, K_\infty,m)$ of $K_{\infty}$ such that
\begin{itemize}
    \item $R^{\Sigma_2}(E/\widetilde{K}_{\infty})$ is a torsion $\Lambda$-module for each $\widetilde{K}_{\infty} \in U$;
    \item $\mu(R^{\Sigma_2}(E/\widetilde{K}_{\infty})) \leq \mu(R^{S_2}(E/K_{\infty}))$ for each $\widetilde{K}_{\infty} \in U$;
    \item $\lambda(R^{\Sigma_2}(E/\widetilde{K}_{\infty})) \leq \lambda(R^{S_2}(E/K_{\infty}))$ for each $\widetilde{K}_{\infty} \in U$ such that $\mu(R^{S_2}(E/\widetilde{K}_{\infty})) = \mu(R^{S_2}(E/K_{\infty}))$.
\end{itemize}
\end{theorem}
\begin{remark}
  If an analogue of  \nameref{conja} holds for $K_\infty$, meaning that $R^S(E/K_\infty)^\vee$ is a finitely generated $\mathbb{Z}_p$-module, then it follows that an analogue of  \nameref{conja} also holds for $R^\Sigma(E/\widetilde{K}_\infty)$ for every $\widetilde{K}_\infty \in \mathcal{U}(K_\infty, m)$ and for every $\Sigma$ containing $S_1$. A similar remark holds for $S_2$.

\end{remark}

\section{An example}\label{example}
Let $p=2$ and $K=\Ff_2(t)$. Let $K_\infty$ be a $\Z_p$-extension constructed using Carlitz module that is ramified only at the prime $\mathfrak{P}=(t)$. Let $E/K$ be an elliptic curve given by the Weierstrass equation:
	
		\begin{equation}
			y^2 + xy = x^3 + (1/t) x^2 + 1,
		\end{equation}	

	 Then $E$ has bad reduction only at the prime $(t)$ of $K$ and is an ordinary elliptic curve. 
Let $S_1$ be the singleton set containing the prime $(t)$ and   $\Sigma$ be a finite set containing $S_1$ . Then, by \cite[Corollary 3.15]{GhoshJhaShekhar}, it follows that $R^{S_1}(E/K_\infty)^\vee$ is a finitely generated $\mathbb{Z}_p$-module. Using Theorem \ref{mainthm:1} (and similarly Theorem \ref{mainthm:2}), we can conclude that there exists a neighbourhood $U = \calU(K_\infty, m)$ (resp. $\calV(K_\infty, m)$) of $K_\infty$ such that for every $\widetilde{K}_\infty \in U$, the module $R^\Sigma(E/\widetilde{K}_\infty)^\vee$ is finitely generated over $\mathbb{Z}_p$. Consequently,  \nameref{conja} of Coates and Sujatha holds for $R^\Sigma(E/\widetilde{K}_\infty)^\vee$ for all $\widetilde{K}_\infty \in U$.
\section*{Acknowledgement}
The  author gratefully acknowledges the support received from the NBHM postdoctoral fellowship.
\iffalse
\section*{Declaration}
During the preparation of this work the author used ChatGPT in order to  enhance the readability of the article. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.
\fi

\bibliographystyle{alpha}
\bibliography{bib}
\end{document}